OpenStudy (anonymous):

Write each equation in slope-intercept form and then find the slope and y-intercept. 2x=y-4

OpenStudy (saifoo.khan):

Slope intercept form => y = mx + b After conversion, "m" is slope. and b is the y-int.

OpenStudy (anonymous):

IDon't Get It. Can You Show Me The Work.?

OpenStudy (anonymous):

y = 2x + 4 so slope = 2 and y-intercept = 4

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Wait. No. 2x=y-4

OpenStudy (saifoo.khan):

2x=y-4 subtract y from both sides, 2x - y = -4 subtract 2x from both sides, -y = -4 - 2x Now multiply the whole equation my -1 as we have to get the "y" in +ve. So, y = 2x + 4 Where 2's the slope and 4 in the intercept.

OpenStudy (anonymous):

How Do ISubtract y from both sides.?

OpenStudy (anonymous):

Wait. So,

OpenStudy (anonymous):

2x=y-4 -2x -2x y=2x+4

OpenStudy (anonymous):

The slope of the line is 2 and the y-intercept is 4. Right.?

OpenStudy (anonymous):

u hv done it wrong!! By your method it should be done as follows 2x = y - 4 Adding 4 to both sides +4 + 4 2x + 4 = y or y = 2x + 4